OpenStudy (anonymous):

solve the Integral (e^(1/x) )/(2x^2)) dx...

myininaya (myininaya):

try the sub on the u=1/x

myininaya (myininaya):

\[-\frac{1}{2} \int\limits_{}^{}e^\frac{1}{x} \cdot \frac{-1}{x^2} dx\]

myininaya (myininaya):

Have you tried it yet?